# ZSMerge: Zero-Shot KV Cache Compression for Memory-Efficient Long-Context LLMs

### v2

@@ 4.1 Experiment Setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA2-7B, LLaMA3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
The baselines for comparison are different KV cache management methods. Specifically, they include:
  • Full KV (FullKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
  • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2). Leverages layer-wise attention score averaging across queries to identify critical tokens, achieving adaptive compression while maintaining task-specific semantics.
  • LESS: Introduces recurrent KV state synthesis through layer-aware dynamic merging. Combines temporal locality exploitation (retaining recent tokens) with spatial compression via parameterized merging operators. Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
  • ZSMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based zero-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
@@ 4.2.2 Throughput Improvement
Macro-Level Analysis
Under diverse workloads, ZSMerge consistently outperformed the baseline and other methods. For moderate batch sizes (batch=4), it achieved 23.4% higher throughput and 18.9% lower latency than full KV caching. At higher batch sizes (batch=8), ZSMerge handled requests that caused OOM errors for other methods, delivering a 124% throughput improvement. For large batch sizes (batch=16), it sustained 2.1× higher throughput and halved latency compared to full KV caching.
In summary, ZSMerge effectively reduces memory usage and maintains high throughput across various workloads, outperforming full KV caching and other baselines. Its dynamic cache compression mechanism ensures scalability and efficient long-context processing without sacrificing performance.
@@ 4.3 Inference Quality Loss
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The experiment compares three KV cache compression methods across two 7B models (LLaMA2-7B and Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. For ZSMerge, the hyperparameters were set to static values based on empirical validation: the residual compensation intensity ratio ($\alpha$) was fixed at $0.6$ to balance performance across tasks, and the historical contribution decay rate ($\lambda$) was set to $0.98$ to ensure stable context propagation while prioritizing recent attention patterns. These values were selected through comparative analysis against other configurations (e.g., $\alpha\in[0.6,0.8]$ and $\lambda=0.98$ showed optimal trade-offs between compression efficiency and generation quality) and required no task-specific training.
Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training-ZSMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets

### v3

@@ 4.1 Experiment Setup
Model Coverage: Our evaluation encompasses diverse modern LLM architectures for broad compatibility. Core experiments use LLaMA2-7B, Falcon-7B, and Mistral-7B-Instruct on NVIDIA A800-80GB GPU. We extend evaluation to LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct, and Yi-6B (Appendix C.4), spanning different attention mechanisms including Multi-Query Attention (MQA).
Benchmark Diversity: We use synthetic datasets for efficiency evaluation and multiple public benchmarks for quality assessment: LongBench (bai2024longbench, ) (21 tasks, 6 categories), XSum (narayanDontGiveMe2018, ) (summarization), InfiniteBench (zhang-etal-2024-bench, ) (100K+ tokens), and GSM-Infinite-8k (zhouGSMInfiniteHowYour2025, ) (mathematical reasoning). Extended results are in Appendix C.
Baselines: We compare ZSMerge against representative KV cache management methods:
  • FullKV: Stores all key-value pairs at every layer, providing the vanilla baseline with maximal memory and compute overhead, but serving as the upper bound for task performance.
  • StreamingLLM (Stream) (xiao2023efficient, ): Retains both early prompt tokens and a fixed sliding window of recent tokens. By designating "attention sinks," it ensures semantic anchors remain available. This static rule is efficient and robust but lacks adaptivity to task-specific token importance.
  • SnapKV (li2024snapkv, ): Performs a one-time cache pruning during prefilling, leveraging early attention distributions to discard less relevant tokens. It eliminates runtime overhead and achieves efficiency, but cannot adapt if token importance shifts during generation.
  • H2O (zhang2024h2o, ): Maintains adaptivity via cumulative attention thresholds that dynamically retain heavy hitters alongside recency-biased tokens. Layer-wise averaging of attention scores captures persistent importance, enabling balanced compression while preserving key semantics.
  • LESS (dong_get_2024, ): Introduces dynamic KV state synthesis through recurrent merging. It combines recency preservation with similarity-based compression, then applies attention rectification to correct distortions introduced by merging. This allows sublinear cache growth without severe loss of contextual coherence.
Extended Baseline Coverage: Beyond the core baselines, our comprehensive evaluation includes additional state-of-the-art methods to provide thorough comparative analysis. These include OmniKV (haoOmniKVDynamicContext2024, ) for attention-guided compression, InfLLM (xiaoInfLLMTrainingFreeLongContext2024, ) for infinite-length modeling, Minference (jiangMInference10Accelerating2024, ) for efficient attention approximation, and FlexPrefill (laiFlexPrefillContextAwareSparse2025, ) for flexible prefilling strategies. Detailed comparisons with these advanced baselines on challenging benchmarks such as InfiniteBench and GSM-Infinite-8k are presented in Appendix C.4, demonstrating ZSMerge’s competitive performance against the latest compression techniques.
@@ 4.2.2 Throughput Improvement
ZSMerge consistently outperformed baselines across diverse workloads, achieving superior throughput while avoiding OOM errors that plagued other methods. Notably, for memory-intensive scenarios (e.g., 13B model, 4096+4096 seq, batch 16), ZSMerge was the only method capable of processing requests (178.2 tokens/s), demonstrating its scalability advantage through dynamic compression.
4.3 Numerical Error Analysis
To further substantiate the role of representation bias correction, we conduct a numerical error analysis that isolates the effect of different compression strategies. In particular, we measure the relative error of attention outputs under varying cache budget constraints, comparing pure eviction-based compression against our residual merging approach. This setup directly reflects the practical conditions under which compression-induced representation shifts occur.
Figure 4: Numerical error analysis comparing eviction vs. residual merging across compression ratios. The log-scale visualization demonstrates that ZSMerge’s residual merging consistently reduces approximation error relative to pure eviction, with benefits amplified under aggressive compression ($\leq 20\%$ cache size).
As shown in Figure 4, under aggressive compression scenarios ($\leq 20\%$ cache size), ZSMerge’s residual merging demonstrates substantial error reduction: 60.5% at 20% cache size, 43.8% at 10%, and 37.4% at 5%. The most dramatic improvement occurs at 50% compression, where residual merging achieves 89.1% error reduction ($2.72\times 10^{-4}$ vs. $2.50\times 10^{-3}$). This empirical validation confirms that the residual compensation mechanism effectively preserves attention distribution coherence, preventing the error propagation characteristic of pure eviction strategies.
The results show that residual merging consistently yields lower approximation error than eviction-based methods, especially under tight memory budgets where aggressive compression is required. This analysis empirically validates the representation-bias-correction hypothesis underlying ZSMerge. Additional validation on the efficiency of residual compensation and its connection to Eq. 8 is provided in Appendix C.2.
@@ 4.4 Inference Quality Loss
4.4.1 Impact of Cache Budget
Table 2: Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
@@ 4.4.1 Impact of Cache Budget
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. Experiments compare three KV-cache compression methods across two 7B-parameter models (LLaMA2-7B and Falcon-7B) under 20%, 10%, and 5% cache budgets. H2O follows its original design, balancing recent tokens and heavy hitters, while LESS uses its recommended merging network trained on the C4 dataset raffel2020exploring . For ZSMerge, budget allocation matches LESS ($B_{r}=2$, with $B_{p}$ and $B_{c}$ sharing the remaining budget), and other hyperparameters are fixed as previously described.
Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. As shown in Table 2, when compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training. ZSMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, its training-free nature enables consistent robustness across different model architectures, avoiding the domain adaptation challenges faced by data-driven methods such as LESS that require training data to tune auxiliary modules.
4.4.2 Generalization across Task Types
Table 3: Results on LongBench benchmark including code completion (CODE), few-shot learning(FSHOT), multi-document QA (MDQA), single-document QA (SDQA), summarization (SUMM), and synthetic reasoning (SYNC) tasks.
Method
CODE
FSHOT
MDQA
SDQA
SUMM
SYNC
LlaMA2-7B
FullKV 65.19 52.32 11.48 17.50 15.15 5.01
Cache size=512
H2O 19.63 7.37 4.48 4.05 3.51 2.16
Stream 61.26 47.63 8.35 11.60 7.07 4.32
SnapKV 63.06 51.51 10.33 15.85 12.28 5.50
ZSMerge 63.02 51.42 10.36 15.67 12.17 5.34
Cache size=1024
H2O 29.12 20.34 5.92 7.47 6.70 2.97
Stream 62.94 49.98 8.16 11.80 7.18 4.45
SnapKV 64.45 52.02 11.30 16.65 13.30 5.09
ZSMerge 64.41 52.06 11.37 16.63 13.23 5.02
Mistra-7B
FullKV 62.10 63.09 37.34 38.94 26.14 66.67
Cache size=512
Stream 57.90 53.68 27.04 23.92 19.80 32.17
SnapKV 60.46 62.24 34.75 36.86 22.18 64.83
ZSMerge 60.48 62.17 34.62 36.83 22.09 64.50
Cache size=1024
Stream 60.13 56.01 27.80 25.10 21.25 34.17
SnapKV 61.60 62.36 35.49 37.30 23.73 66.50
ZSMerge 61.59 62.42 35.43 37.31 23.64 66.50
We evaluate the effectiveness of ZSMerge on the LongBench benchmark, which covers a diverse set of task types, including code completion, few-shot learning, multi-document QA, single-document QA, summarization, and synthetic reasoning tasks. Experiments are conducted on two backbone models: LLaMA2-7B and Mistral-7B, with cache size constraints set to 512 and 1024. For the Mistral-7B model, the H2O baseline encountered out-of-memory (OOM) errors and could not be included in the comparison. More details about the experimental setup and validation results on specific datasets are provided in Appendix C.
Overall, the results in Table 3 show that ZSMerge consistently matches or closely tracks the performance of FullKV across most task types, despite operating under strict cache constraints. Compared to Stream and H2O, ZSMerge demonstrates substantially higher accuracy, especially in complex and memory-intensive tasks like multi-document QA and summarization. Interestingly, while SnapKV benefits from a one-time pruning strategy during the prefilling phase which allows it to retain a larger cache during decoding, ZSMerge still achieves comparable or even slightly better performance in several tasks, thanks to its adaptive two-phase compression and positional-aware merging strategy.
These findings highlight ZSMerge’s strong generalization and robustness across heterogeneous workloads, making it a promising solution for real-world deployment in memory-constrained scenarios. Additional results with richer benchmarks and broader baseline comparisons are provided in Appendix C.4.